Find the coordinates of the points on the curve 𝑦 = 3π‘₯3 βˆ’ 2π‘₯2 βˆ’ 12π‘₯ + 2 where the normal is parallel to the line 𝑦 = βˆ’π‘₯ + 1.


1
Expert's answer
2022-03-03T18:03:59-0500
"y=3x^3-2x^2-12x+2"

"y'=9x^2-4x-12"

The slope of the normal is


"slope=-\dfrac{1}{f'(x_0)}=-1"

Then


"9x_0^2-4x_0-12=1"

"9x_0^2-4x_0-13=0"

"x_0=\dfrac{13}{9}\ or\ x_0=-1"

"x_0=\dfrac{13}{9}"


"y(\dfrac{13}{9})=3(\dfrac{13}{9})^3-2(\dfrac{13}{9})^2-12(\dfrac{13}{9})+2=-\dfrac{2543}{243}"

"x_0=-1"


"y(-1)=3(-1)^3-2(-1)^2-12(-1)+2=9"

"Point(\dfrac{13}{9}, -\dfrac{2543}{243}), Point(-1, 9)"